40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Nov 3 08:07:52 2016, Steve, Update, General, power glitch powerGlitch.png
    Reply  Thu Nov 3 11:33:24 2016, gautam, Update, General, power glitch - recovery 
Message ID: 12594     Entry time: Thu Nov 3 11:33:24 2016     In reply to: 12593
Author: gautam 
Type: Update 
Category: General 
Subject: power glitch - recovery 

I did the following:

  • Hard reboots for fb, megatron, and all the frontends, in that order
  • Checked time on all FEs, ran sudo ntpdate -b -s -u pool.ntp.org where necessary
  • Restarted all realtime models
  • Restarted monit on all FEs
  • Reset Marconi to nominal settings, fCarrier=11.066209MHz, +13dBm amplitude
  • In the control room, restarted the projector and set up the usual StripTool traces
  • Realigned PMC
  • Slow machines did not need any touchups - interestingly, ITMX did not get stuck during this power glitch!

There was a regular beat coming from the speakers. After muting all the channels on the mixer and pulling the 3.5mm cable out, the sound persisted. It now looks like the mixer is broken sad

     ProFX8v2

 

ELOG V3.1.3-